This is a calculus problem.
Consider two integers x and y such that x−y=12. What value of y will minimize x^2+y^2?

Answer by richwmiller(17219)

min{z = 2 (x^2-12 x+72)} = 72 at x = 6
parabola
focus | (6, 577/8)~~(6, 72.125)
vertex | (6, 72)
semi-axis length | 1/8 = 0.125
focal parameter | 1/4
eccentricity | 1
directrix | z = 575/8